# UniAudio 2.0: A Unified Audio Language Model with Text-Aligned Factorized Audio Tokenization

### v2

@@ 1 Introduction
Figure 1: The overview of the proposed UniAudio 2.0.
@@
Figure 2: An overview of the framework of the proposed ReasoningCodec. Semantic Decoder denotes the semantic feature decoder, which consists of several convolutional layers.
@@ 2.1 Audio Language Models
Recent years have witnessed rapid progress in audio language models that aim to bridge audio and text through multimodal learning (Xu et al., 2025; Zhang et al., 2025; Borsos et al., 2023; Ding et al., 2025). Existing approaches largely fall into two distinct paradigms, depending on how audio representations are integrated with language models. The first paradigm focuses on audio understanding by coupling continuous audio representations with pre-trained language models (Chu et al., 2024; Tang et al., 2024; Xu et al., 2025). In this line of work, audio signals are encoded into continuous features using pre-trained audio encoders (Radford et al., 2023; Chen et al., 2022; Hsu et al., 2021), which are then aligned with textual representations to support perception and reasoning tasks. The second paradigm formulates audio modeling as a discrete sequence prediction problem, drawing inspiration from autoregressive text language models (Borsos et al., 2023; Wang et al., 2023; Kharitonov et al., 2023; Yang et al., 2023c). In this setting, raw audio is first converted into a sequence of discrete tokens using an audio tokenizer or codec (Zeghidour et al., 2021; Défossez et al., 2022; Kumar et al., 2023; Zhang et al., 2023; Yang et al., 2023b), and an autoregressive transformer is trained to model and generate audio token sequences. This paradigm has been successfully applied to text-to-speech (Wang et al., 2023; Kharitonov et al., 2023), music generation (Copet et al., 2023; Agostinelli et al., 2023), speech-to-speech dialogue (Défossez et al., 2024; Zeng et al., 2024; Ding et al., 2025; Nguyen et al., 2025), and unified multi-task audio generation models (Yang et al., 2023c; Liu et al., 2025a; Wang et al., 2024; Vyas et al., 2023). Recently, more work has focused on building unified speech understanding and generation models under the LM paradigm, such as OpusLM (Tian et al., 2025b), DualSpeechLM (Wang et al., 2025c), and Ming-UniAudio (Yan et al., 2025). In this work, we focus on building unified audio understanding and generation models that can understand and generate text, speech, sound, and music.
@@ 3 ReasoningCodec
A core challenge for unified audio models is to design an intermediate audio representation that is simultaneously LM-friendly for autoregressive modeling and information-preserving for LM-based understanding. Prior work (Borsos et al., 2023; Yang et al., 2025a) shows that semantic discrete tokens are generally easier to model than purely acoustic tokens in autoregressive generation. However, directly using discrete semantic tokens for understanding remains non-trivial: vector quantization introduces information loss, which often degrades performance on comprehension tasks (see Table 3). To bridge this gap, we propose ReasoningCodec, a factorized audio tokenizer that decomposes an audio waveform into two complementary token streams: (i) Reasoning tokens, which encode text-aligned, high-level perceptual analyses and planning representations, optimized to match the inductive biases of text LLMs for efficient understanding and generation. Here, reasoning denotes grounded perceptual inference over acoustic cues, rather than explicit multi-step chain-of-thought reasoning in text-based LLMs (Guo et al., 2025). (ii) Multi-level reconstruction tokens that capture semantic content and fine-grained acoustics for high-fidelity waveform reconstruction and LM-based autoregressive generation. Formally, given an audio $x$, ReasoningCodec produces $r=\mathcal{T}_{r}(x)$ and $s=\mathcal{T}_{s}(x\mid r)$, where $r$ denotes reasoning tokens and $s$ denotes multi-level reconstruction tokens. $\mathcal{T}_{r}$ and $\mathcal{T}_{s}$ denote the reasoning branch and reconstruction branch, respectively. Note that the waveform is reconstructed only from $\hat{x}=\mathcal{D}(s)$. Figure 2 shows the framework of ReasoningCodec. In the following, we present the details of the reasoning branch, the reconstruction branch, and the training procedure of ReasoningCodec.
@@ 4.4 Data and Training Strategy
UniAudio 2.0 is trained with a diverse collection of text and audio data under a multi-task, multi-stage training paradigm. Our data construction and training strategy are designed to (i) preserve the strong textual capability inherited from pre-trained LLMs, (ii) progressively inject audio understanding and generation abilities, and (iii) improve generalization to unseen tasks.
Multi-task Data Construction We organize the training corpus into several complementary data types, each corresponding to a class of tasks supported by UniAudio 2.0. The training data include (1) text-only data, (2) audio-only data, (3) speech-transcription paired data, (4) speech-caption-transcription paired data, (5) audio/music-caption paired data, (6) lyric-song paired data, and (7) auditory sentences constructed using our proposed task-construction strategy. We provide details of the data and tasks in Appendix B.2.
Multi-stage Training Strategy
@@
XCodec 9.22 16.31 21.73 26.36 18.40 11.77 23.60 34.47 42.3 28.5
HiggCodec 9.12 29.49 51.32 65.92 38.46 12.83 21.34 57.82 88.1 45.02
DAC Codec 13.74 66.06 116.54 163.89 90.06 25.67 68.09 99.66 133.9 81.34
Reason-only 3.68 7.73 15.88 18.99 11.57 3.81 6.97 7.26 8.20 6.56
Reconstruction-only 8.69 13.99 19.75 75.02 29.36 9.19 22.94 15.94 59.3 26.33
@@
We adopt a four-stage training strategy that progressively integrates audio understanding and generation capabilities into the unified autoregressive model, including: (1) Stage 1: Audio understanding warm-up, (2) Stage 2: Audio generation warm-up; (3) Stage 3: Audio-text pre-training; (4) Stage 4: Audio-text mid-training. Figure 5 presents an overview of the four-stage training.
Stage 1: Audio understanding warm-up In the first stage, we focus on initializing the audio understanding experts. We train the model using a subset of audio understanding tasks, while freezing all other components. To encourage the understanding experts to encode rich semantic information, we introduce an auxiliary semantic distillation objective. Following the training of ReasoningCodec 3.2, a lightweight decoder is attached to reconstruct semantic features extracted from frozen WavLM and music SSL. The overall objective includes reconstruction loss and text lm loss. After the training, we will discard the decoder.
Stage 2: Audio generation warm-up We train the audio generation expert and the local audio decoder. We train the model on a subset of audio generation tasks while keeping the understanding and cross-modal experts fixed.
Stage 3: Audio-text pre-training We jointly update all model parameters using a mixture of audio understanding tasks, audio generation tasks, text-only data, and audio-only data. This stage aligns the two modalities under the autoregressive objective. In this stage, the largest context is 1024.
Stage 4: Audio-text mid-training In the final stage, we aim to extend the effective context length and enhance generalization to unseen tasks. We continue training on a subset of the pre-training data in stage 3, augmented with our constructed auditory sentence data. This stage encourages the model to reason over longer and more complex audio-text sequences and improves robustness across diverse task settings. In this stage, the largest context is 2048.
We provide more details about each training stage in the Appendix B.3.
@@ 4.5 Connection to the UniAudio Series
Similar to UniAudio and UniAudio 1.5, UniAudio 2.0 aims to develop a general-purpose audio foundation model that supports speech, sound, and music understanding and generation within a unified framework. While earlier systems primarily focused on multi-task learning and in-context adaptation, UniAudio 2.0 further emphasizes representation-level alignment between audio and language, enabling more scalable and transferable modeling.
@@ 5.2 Evaluation Metrics
Audio Tokenizer Evaluation We evaluate the performance of audio tokenizers in (1) audio reconstruction, (2) LLM-based audio understanding, and (3) LLM-based audio generation.
@@
Audio Reconstruction For speech reconstruction, we use DNS-MOS, UT-MOS, PESQ, STOI (Short-time Objective Intelligibility), and VISQOL. For sound and music data evaluation, we follow (Boson AI, 2025), using VISQOL (audio version) and AudioBox aesthetics score. Furthermore, following (Kumar et al., 2023), the MUSHRA subjective test is conducted for speech, sound, and music. Refer to Appendix A.3.4 for more details.
@@
LLM-based audio understanding tasks To validate whether the discrete audio tokenizer is suitable as an intermediate representation for LLM-based audio understanding tasks, we follow the setting of Qwen-Audio (Chu et al., 2024; Tang et al., 2024) to conduct multi-task training, including ASR (Panayotov et al., 2015), emotion recognition (21), audio classification (Mesaros et al., 2017), and music classification (Tzanetakis et al., 2001). When we train the model on these downstream tasks, only the adapter and LoRA (Hu et al., 2021) are updated.
@@
LLM-based audio generation tasks To validate whether the proposed audio tokenizer is suitable for AR modeling, we follow (Yang et al., 2025a; Wang et al., 2025b), using PPL and token prediction accuracy as the metric.
Audio understanding and generation tasks evaluation In this study, we focus on multiple audio-related understanding and generation tasks. To better demonstrate the capacity of UniAudio 2.0, we design three levels of evaluation: (1) seen-task performance; (2) few-shot evaluation; and (3) zero-shot evaluation. We list the details of these tasks in Appendix B.4.
@@
Seen tasks These tasks are seen during the pre-training stage, including TTS, Instruct TTS, ASR, audio generation/caption, music generation/caption, song generation, and lyric recognition. For these tasks, we follow the commonly used benchmark and evaluation metrics.
@@
Few-shot tasks We also design several few-shot tasks, including speech denoising, voice conversion, emotion recognition, and audio classification. For each task, we consider one-shot and two-shot settings. Note that MiMo-Audio (Zhang et al., 2025) also introduces a few-shot setting, but our context length is 2048, which is not suitable for using the same setup as MiMo-Audio.
@@
Zero-shot tasks Lastly, we expect UniAudio 2.0 to exhibit zero-shot generalization to unseen tasks. To this end, we evaluate it on several tasks, including text question answering, speech-to-speech conversation, dysarthric speech recognition, speech-to-sound generation, and audio-prompted instruction-following TTS. For these tasks, we do not explicitly include any corresponding task-specific training data; therefore, we treat them as zero-shot evaluations.
5.3 The Performance of Audio Tokenizer
@@
Reconstruction Performance Table 1 presents reconstruction performance on speech, sound, and music evaluation sets. We compare against previous SOTA universal audio codecs, including DAC (Kumar et al., 2023), Encodec (Défossez et al., 2022), Higg-AudioCodec (Boson AI, 2025), X-Codec (Ye et al., 2025), and ALMTokenizer (Yang et al., 2025a). We observe that the proposed ReasoningCodec achieves strong reconstruction quality across different audio modalities. In Appendix A.3.4, we conduct subjective evaluations of different audio tokenizers. As Table 14 shows, at the same token rate, ReasoningCodec achieves better reconstruction performance.
@@
Token Modeling Performance We explore whether ReasoningCodec is suitable for LLM-based audio generation and understanding tasks. For the generation experiments, we use the same LLM backbone, training data, and evaluation data, but choose different audio tokenizers. As Table 2 shows, we observe that (1) semantically enhanced audio tokenizers (e.g., XCodec) perform better than purely acoustic tokenizers (e.g., DAC); (2) our proposed ReasoningCodec achieves strong performance in PPL and token prediction accuracy; (3) reasoning tokens are easier for an LM to model, and combining reasoning tokens with reconstruction tokens significantly improves the prediction accuracy of reconstruction tokens. These results further demonstrate the effectiveness of ReasoningCodec.
@@
Table 3 shows the results of audio understanding tasks. We can see that ReasoningCodec achieves the best understanding performance among discrete audio tokenizers. Furthermore, its performance is close to that of a continuous audio tokenizer (Whisper) on multiple audio understanding tasks. Figure 3 shows the training loss for each tokenizer; the training loss of reasoning tokens decreases very quickly: these phenomenon further validate the effectiveness of ReasoningCodec. In Section 5.4, we show that UniAudio 2.0 built on ReasoningCodec achieves comparable performance with other audio understanding competitors. In Appendix A.3.5, we present a more detailed analysis of why introducing reasoning tokens is effective.

### v3

@@ 1 Introduction
Figure 1: Overview of the proposed UniAudio 2.0 framework. UniAudio 2.0 adopts a unified autoregressive architecture over text, reasoning, and reconstruction tokens, where reasoning tokens capture high-level, text-grounded semantics and reconstruction tokens preserve fine-grained acoustic details. The model integrates audio understanding, cross-modal, and audio generation experts to support unified audio understanding and generation.
@@
Figure 2: Figure X. Overview of the proposed ReasoningCodec framework. ReasoningCodec adopts a dual-branch architecture, consisting of a reasoning branch and a reconstruction branch, which are coupled through FiLM-based modulation. The semantic decoder consists of several convolutional layers.
@@ 2.1 Audio Language Models
Recent years have witnessed rapid progress in audio language models that aim to bridge audio and text through multimodal learning (Xu et al., 2025; Zhang et al., 2025; Borsos et al., 2023; Ding et al., 2025). Existing approaches largely fall into two distinct paradigms, depending on how audio representations are integrated with language models. The first paradigm focuses on audio understanding by coupling continuous audio representations with pre-trained language models (Chu et al., 2024; Tang et al., 2024; Xu et al., 2025). In this line of work, audio signals are encoded into continuous features using pre-trained audio encoders (Radford et al., 2023; Chen et al., 2022; Hsu et al., 2021), which are then aligned with textual representations to support perception and reasoning tasks. The second paradigm formulates audio modeling as a discrete sequence prediction problem, drawing inspiration from autoregressive text language models (Borsos et al., 2023; Wang et al., 2023; Kharitonov et al., 2023; Yang et al., 2023c). In this setting, raw audio is first converted into a sequence of discrete tokens using an audio tokenizer or codec (Zeghidour et al., 2021; Défossez et al., 2022; Kumar et al., 2023; Zhang et al., 2023b; Yang et al., 2023b), and an autoregressive transformer is trained to model and generate audio token sequences. This paradigm has been successfully applied to text-to-speech (Wang et al., 2023; Kharitonov et al., 2023), music generation (Copet et al., 2023; Agostinelli et al., 2023), speech-to-speech dialogue (Défossez et al., 2024; Zeng et al., 2024; Ding et al., 2025; Nguyen et al., 2025), and unified multi-task audio generation models (Yang et al., 2023c; Liu et al., 2025a; Wang et al., 2024; Vyas et al., 2023). Recently, more work has focused on building unified speech understanding and generation models under the LM paradigm, such as OpusLM (Tian et al., 2025b), DualSpeechLM (Wang et al., 2025c), and Ming-UniAudio (Yan et al., 2025). In this work, we focus on building unified audio understanding and generation models that can understand and generate text, speech, sound, and music.
@@ 3 ReasoningCodec
A core challenge for unified audio models is to design an intermediate audio representation that is simultaneously LM-friendly for autoregressive modeling and information-preserving for LM-based understanding. Prior work (Borsos et al., 2023; Yang et al., 2025a) shows that semantic discrete tokens are generally easier to model than purely acoustic tokens in autoregressive generation. However, directly using discrete semantic tokens for understanding remains non-trivial: vector quantization introduces information loss, which often degrades performance on comprehension tasks (see Table 3). To bridge this gap, we propose ReasoningCodec, a factorized audio tokenizer that decomposes an audio waveform into two complementary token streams: (i) Reasoning tokens  [^2] Note that reasoning tokens differ from conventional semantic tokens. While prior semantic tokenizers typically capture both text-level semantics and non-textual paralinguistic information such as prosody, our reasoning tokens are deliberately designed to capture only text-level semantic information. As a result, reasoning tokens contain less information than traditional semantic tokens and focus on high-level, text aligned semantics., which encode text-aligned, high-level perceptual analyses and planning representations, optimized to match the inductive biases of text LLMs for efficient understanding and generation. Here, reasoning denotes grounded perceptual inference over acoustic cues, rather than explicit multi-step chain-of-thought reasoning in text-based LLMs (Guo et al., 2025). (ii) Multi-level reconstruction tokens that capture semantic content and fine-grained acoustics for high-fidelity waveform reconstruction and LM-based autoregressive generation. Formally, given an audio $x$, ReasoningCodec produces $r=\mathcal{T}_{r}(x)$ and $s=\mathcal{T}_{s}(x\mid r)$, where $r$ denotes reasoning tokens and $s$ denotes multi-level reconstruction tokens. $\mathcal{T}_{r}$ and $\mathcal{T}_{s}$ denote the reasoning branch and reconstruction branch, respectively. Note that the waveform is reconstructed only from $\hat{x}=\mathcal{D}(s)$. Figure 2 shows the framework of ReasoningCodec. In the following, we present the details of the reasoning branch, the reconstruction branch, and the training procedure of ReasoningCodec.
@@ 4.4 Data and Training Strategy
UniAudio 2.0 is trained on a diverse collection of text and audio data under a multi-task, multi-stage training paradigm. Our data construction and training strategy are designed to: (i) preserve the strong textual capability inherited from pre-trained LLMs, (ii) progressively inject audio understanding and generation abilities, and (iii) improve generalization to unseen tasks.
Multi-task Data Construction. We organize the training corpus into several complementary data types, each corresponding to a class of tasks supported by UniAudio 2.0. The training data include: (1) text-only data, (2) audio-only data, (3) speech–transcription paired data, (4) speech–caption–transcription paired data, (5) audio/music–caption paired data, (6) lyric–song paired data, and (7) auditory sentences constructed using our proposed task-construction strategy. Detailed descriptions of the data and tasks are provided in Appendix B.2.
Multi-stage Training Strategy. We adopt a four-stage training strategy to progressively integrate audio understanding and generation capabilities into the unified autoregressive model, including: (1) Stage 1: Audio understanding warm-up, (2) Stage 2: Audio generation warm-up, (3) Stage 3: Audio–text pre-training, and (4) Stage 4: Audio–text mid-training. Figure 5 provides an overview of the four training stages.
Stage 1: Audio Understanding Warm-up. In the first stage, we focus on initializing the audio understanding experts. The model is trained on a subset of audio understanding tasks while all other components are frozen. To encourage the understanding experts to encode rich semantic information, we introduce an auxiliary semantic distillation objective. Following the training of ReasoningCodec (Section 3.2), a lightweight decoder is attached to reconstruct semantic features extracted from frozen WavLM and music SSL models. The overall objective consists of a reconstruction loss and a language modeling loss. After training, the auxiliary decoder is discarded.
Stage 2: Audio Generation Warm-up. In this stage, we train the audio generation expert and the local audio decoder. The model is optimized on a subset of audio generation tasks, while the understanding and cross-modal experts remain fixed.
Stage 3: Audio–Text Pre-training. We jointly update all model parameters using a mixture of audio understanding tasks, audio generation tasks, text-only data, and audio-only data. This stage aligns the two modalities under a unified autoregressive objective. The maximum context length in this stage is 1024.
Stage 4: Audio–Text Mid-training. In the final stage, we aim to extend the effective context length and enhance generalization to unseen tasks. We continue training on a subset of the pre-training data from Stage 3, augmented with the constructed auditory sentence data. This stage encourages the model to reason over longer and more complex audio–text sequences and improves robustness across diverse task settings. The maximum context length in this stage is 2048.
Additional details of each training stage are provided in Appendix B.3.
@@
XCodec (Ye et al., 2025) 9.22 16.31 21.73 26.36 18.40 11.77 23.60 34.47 42.3 28.5
HiggCodec (Boson AI, 2025) 9.12 29.49 51.32 65.92 38.46 12.83 21.34 57.82 88.1 45.02
DAC Codec (Kumar et al., 2023) 13.74 66.06 116.54 163.89 90.06 25.67 68.09 99.66 133.9 81.34
Reason-only (Ours) 3.68 7.73 15.88 18.99 11.57 3.81 6.97 7.26 8.20 6.56
Reconstruction-only (Ours) 8.69 13.99 19.75 75.02 29.36 9.19 22.94 15.94 59.3 26.33
Reason + Reconstruction (Ours) 6.92 11.37 19.42 55.36 23.77 7.02 15.55 15.23 42.5 20.1
@@
MimiCodec (8 RVQ) (Défossez et al., 2024) 12.5/100 2048 86.7 $\pm$ 2.1 - -
ReasoningCodec (Ours) 12.5/100 1024 90.5 $\pm$ 2.8 80.8 $\pm$ 2.0 86.6 $\pm$ 2.3
@@ 4.5 Connection to the UniAudio Series
Similar to UniAudio and UniAudio 1.5, UniAudio 2.0 aims to develop a general-purpose audio foundation model that supports speech, sound, and music understanding and generation within a unified framework. While earlier systems primarily focused on multi-task learning and in-context adaptation, UniAudio 2.0 further emphasizes representation-level alignment between audio and text, enabling more scalable and transferable modeling.
@@ 5.2 Evaluation Metrics
Audio Tokenizer Evaluation. We evaluate audio tokenizers from three perspectives: (1) audio reconstruction, (2) LLM-based audio understanding, and (3) LLM-based audio generation.
@@
Audio Reconstruction. For speech reconstruction, we adopt DNS-MOS, UT-MOS, PESQ, STOI (Short-Time Objective Intelligibility), and VISQOL. For sound and music evaluation, we follow (Boson AI, 2025) and use VISQOL (audio version) and the AudioBox aesthetics score. In addition, following (Kumar et al., 2023), we conduct MUSHRA subjective evaluations on speech, sound, and music. As shown in Table 4, ReasoningCodec achieves consistently strong reconstruction performance across all modalities.
@@
LLM-based Audio Understanding Tasks. To assess whether discrete audio tokenizers are suitable as intermediate representations for LLM-based audio understanding, we follow the settings of Qwen-Audio (Chu et al., 2024; Tang et al., 2024) and conduct multi-task training on ASR (Panayotov et al., 2015), emotion recognition (21), audio classification (Mesaros et al., 2017), and music classification (Tzanetakis et al., 2001). During downstream task training, only the adapter and LoRA modules (Hu et al., 2021) are updated.
@@
LLM-based Audio Generation Tasks. To evaluate whether the proposed tokenizer is suitable for autoregressive modeling, we follow (Yang et al., 2025a; Wang et al., 2025b) and adopt perplexity (PPL) and token prediction accuracy as evaluation metrics.
Audio Understanding and Generation Tasks Evaluation. We evaluate UniAudio 2.0 on a diverse set of audio-related understanding and generation tasks. To comprehensively assess its capabilities, we consider three evaluation settings: (1) seen-task evaluation, (2) few-shot evaluation, and (3) zero-shot evaluation. Detailed task descriptions are provided in Appendix B.4.
@@
Seen Tasks. These tasks are observed during pre-training, including TTS, instructed TTS, ASR, audio generation and captioning, music generation and captioning, song generation, and lyric recognition. For these tasks, we follow commonly adopted benchmarks and evaluation metrics.
@@
Few-shot Tasks. We design several few-shot tasks, including speech denoising, voice conversion, emotion recognition, and audio classification. For each task, we consider both one-shot and two-shot settings. We note that MiMo-Audio (Zhang et al., 2025) also introduces a few-shot evaluation protocol; however, due to our context length limitation of 2048, we are unable to adopt the same experimental setup.
@@
Zero-shot Tasks. Finally, we evaluate UniAudio 2.0 on a set of unseen tasks to assess its zero-shot generalization ability, including text question answering, speech-to-speech conversation, dysarthric speech recognition, speech-to-sound generation, and audio-prompted instruction-following TTS. For these tasks, we do not include any corresponding task-specific training data, and therefore treat them as zero-shot evaluations.
5.3 Performance of the Audio Tokenizer
@@
Reconstruction Performance. Table 1 presents the reconstruction performance on speech, sound, and music evaluation sets. We compare our method with previous state-of-the-art universal audio codecs, including DAC (Kumar et al., 2023), Encodec (Défossez et al., 2022), Higg-AudioCodec (Boson AI, 2025), X-Codec (Ye et al., 2025), and ALMTokenizer (Yang et al., 2025a). We observe that the proposed ReasoningCodec achieves strong reconstruction quality across different audio modalities. In Table 4,, we further conduct subjective evaluations of different audio tokenizers. We can see that at the same token rate (TPS), ReasoningCodec achieves consistently better reconstruction performance.
@@
Token Modeling Performance. We investigate whether ReasoningCodec is suitable for LLM-based audio generation and understanding tasks. For generation experiments, we use the same LLM backbone, training data, and evaluation data, while varying only the audio tokenizer. As shown in Table 2, we observe that: (1) semantically enhanced audio tokenizers (e.g., X-Codec (Ye et al., 2025)) perform better than purely acoustic tokenizers (e.g., DAC-Codec (Kumar et al., 2023)); (2) our proposed ReasoningCodec achieves strong performance in terms of perplexity and token prediction accuracy; and (3) reasoning tokens are easier for language models to capture, and combining reasoning tokens with reconstruction tokens significantly improves the prediction accuracy of reconstruction tokens. These results further demonstrate the effectiveness of ReasoningCodec.
@@
Table 3 reports the results on audio understanding tasks. The results show that ReasoningCodec achieves the best understanding performance among discrete audio tokenizers, and its performance is close to that of a continuous tokenizer (Whisper) on multiple benchmarks. Figure 3 illustrates the training loss for different tokenizers. Notably, the training loss of reasoning tokens decreases rapidly, and these phenomena further validate the effectiveness of ReasoningCodec. In Section 5.4, we further show that UniAudio 2.0 built on ReasoningCodec achieves competitive performance compared with other audio understanding systems. In Appendix A.3.5, we provide a more detailed analysis of why introducing reasoning tokens is effective.